OpenStudy (anonymous):

Calculate the midpoint between (2, -5) and (4, -7). (1, 1) (3, -6) (-1, -1) (6, -12)

OpenStudy (anonymous):

@JMark

OpenStudy (michele_laino):

coordinates of middle point, are: \[x _{M}=\frac{ x _{1}+x _{2} }{ 2 },y _{M}=\frac{ y _{1}+y _{2} }{ 2 }\] where (x_1,y_1) and (x_2,y_2) are your points

OpenStudy (michele_laino):

@kittymeow101 please substitute your coordinates into formulas above

OpenStudy (anonymous):

i really dont understand this

OpenStudy (michele_laino):

why? for example: \[x _{M}=\frac{ 2+4 }{ 2 }=3\] please do the same, to evaluate y_M

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

\[\frac{ 2-5 }{ 2 }=\frac{ 3 }{ 2}\]

OpenStudy (anonymous):

im sure that wrong...

OpenStudy (michele_laino):

\[y _{M}=\frac{ -5-7 }{ 2 }=-6\]

OpenStudy (anonymous):

oh now i get it

OpenStudy (michele_laino):

so option B. M=(3,-6)

OpenStudy (anonymous):

ok ty so much i have a few more Q's ill close this and give you medals if you can please help

OpenStudy (michele_laino):

ok!